# Block encoding of matrix product operators

### v2

@@ Appendix A Signal processing circuit
$\Pi_{\phi}=\begin{pmatrix}e^{-i\phi}&&&&\\ &e^{i\phi}&&&\\ &&e^{i\phi}&&\\ &&&\ddots&\\ &&&&e^{i\phi}\end{pmatrix}.$ (29aq)
@@
Let us consider the $q$-th operator in the circuit (29ar). This operator has $q-1$ controls and a rotation phase of $\frac{-\phi}{2^{n-q}}$. Its matrix expression, considering the tensor multiplication with $n-q$ identities as well, is:
$\begin{pmatrix}\overbrace{\begin{bmatrix}e^{-i\frac{\phi}{2^{n-q}}}&\\ &\ddots\end{bmatrix}}^{2^{n-q}}&&\\ &\overbrace{\begin{bmatrix}e^{i\frac{\phi}{2^{n-q}}}&\\ &\ddots\end{bmatrix}}^{2^{n-q}}&\\ &&\overbrace{\begin{bmatrix}1&\\ &\ddots\\ \end{bmatrix}}^{2^{n-q}\left(2^{q}-2\right)}\end{pmatrix}$ (29as)
By multiplying all the matrices (29as) with $q$ from 1 to $n$, we get then the following:
$\begin{pmatrix}\displaystyle\prod_{q=1}^{n}e^{-i\frac{\phi}{2^{n-q}}}&&&&&\\ &\displaystyle e^{i\phi}\prod_{q=1}^{n-1}e^{-i\frac{\phi}{2^{n-q}}}&&&&\\ &&\overbrace{\begin{bmatrix}\displaystyle e^{i\frac{\phi}{2}}\prod_{q=1}^{n-2}e^{-i\frac{\phi}{2^{n-q}}}&\\[10.00002pt] &\ddots\end{bmatrix}}^{2}&&&\\ &&&\overbrace{\begin{bmatrix}\displaystyle e^{i\frac{\phi}{4}}\prod_{q=1}^{n-3}e^{-i\frac{\phi}{2^{n-q}}}&\\[10.00002pt] &\ddots\end{bmatrix}}^{4}&&\\ &&&&\ddots&\\ &&&&&\overbrace{\begin{bmatrix}e^{i\frac{\phi}{2^{n-1}}}&\\ &\ddots\end{bmatrix}}^{2^{n-1}}\end{pmatrix}$ (29at)
@@
$=e^{i\phi\left(\frac{1}{2^{n-1}}-2\right)}$ (29av)
By multiplying both equations (A) and (A) by a global phase $e^{i\phi\left(1+\frac{1}{2^{n-1}}\right)}$ we get then the reference matrix from equation (29aq), which proves that the two circuits are equivalent up to a global phase.

### v3

@@ Appendix A Signal processing circuit
$\Pi_{\phi}=\begin{pmatrix}e^{-i\phi}&&&&\\ &e^{i\phi}&&&\\ &&e^{i\phi}&&\\ &&&\ddots&\\ &&&&e^{i\phi}\end{pmatrix}.$ (48)
@@
[Uncaptioned image] (49)
Let us consider the $q$-th operator in the circuit (49). This operator has $q-1$ controls and a rotation phase of $\frac{-\phi}{2^{n-q}}$. Its matrix expression, considering the tensor multiplication with $n-q$ identities as well, is:
$\begin{pmatrix}\overbrace{\begin{bmatrix}e^{-i\frac{\phi}{2^{n-q}}}&\\ &\ddots\end{bmatrix}}^{2^{n-q}}&&\\ &\overbrace{\begin{bmatrix}e^{i\frac{\phi}{2^{n-q}}}&\\ &\ddots\end{bmatrix}}^{2^{n-q}}&\\ &&\overbrace{\begin{bmatrix}1&\\ &\ddots\\ \end{bmatrix}}^{2^{n-q}\left(2^{q}-2\right)}\end{pmatrix}$ (50)
By multiplying all the matrices (50) with $q$ from 1 to $n$, we get then the following:
$\begin{pmatrix}\displaystyle\prod_{q=1}^{n}e^{-i\frac{\phi}{2^{n-q}}}&&&&&\\ &\displaystyle e^{i\phi}\prod_{q=1}^{n-1}e^{-i\frac{\phi}{2^{n-q}}}&&&&\\ &&\overbrace{\begin{bmatrix}\displaystyle e^{i\frac{\phi}{2}}\prod_{q=1}^{n-2}e^{-i\frac{\phi}{2^{n-q}}}&\\[10.00002pt] &\ddots\end{bmatrix}}^{2}&&&\\ &&&\overbrace{\begin{bmatrix}\displaystyle e^{i\frac{\phi}{4}}\prod_{q=1}^{n-3}e^{-i\frac{\phi}{2^{n-q}}}&\\[10.00002pt] &\ddots\end{bmatrix}}^{4}&&\\ &&&&\ddots&\\ &&&&&\overbrace{\begin{bmatrix}e^{i\frac{\phi}{2^{n-1}}}&\\ &\ddots\end{bmatrix}}^{2^{n-1}}\end{pmatrix}$ (51)
@@
$=e^{i\phi\left(\frac{1}{2^{n-1}}-2\right)}$ (53)
By multiplying both equations (A) and (A) by a global phase $e^{i\phi\left(1+\frac{1}{2^{n-1}}\right)}$ we get then the reference matrix from equation (48), which proves that the two circuits are equivalent up to a global phase.
Appendix B Post-selection success probability for systems of interest
Figure 9: Success probabilities for LCU and MPO block encodings, given by Eq. (24) and Eq. (III). The physical systems are, in order, the Ising, Heisenberg, exponentially decaying XY (with $\gamma=0.3$), and Fermi-Hubbard models. The initial state |$\ket{\psi_{\text{in}}}$ has been chosen in all cases equal to |$\ket{+}^{\otimes n}$. A rescaling of the Hamiltonian parameters can drastically change the probability of success for the MPO block encoding, while it does not have any impact on LCU.
In this section, we provide a numerical analysis concerning the probability of success for both MPO and LCU block encodings (without considering the entire QET circuit). The two methods are summarized respectively in Eq. (III) and Eq. (24), where the rescaling factors $N_{\text{MPO}}$ and $N_{\text{LCU}}$ are defined by Eq. (11) and Eq. (23). The post-selection success probability scales then exponentially in $L$ for the MPO block encoding and linearly in $M$ for the LCU method.
We remark that the number of unitaries $M$ of the LCU decomposition can scale exponentially in $L$ in the worst-case scenario but also linearly for many systems of interest, such as Ising, Heisenberg, etc. This means that our method, in the asymptotic limit, presents a much lower success probability compared to LCU in many circumstances.
However, in section VI, and more specifically in Fig. 6, we discussed how rescaling the Hamiltonian might help to reduce the normalization factor so that $N_{\text{MPO}}\rightarrow 1$. This trick doesn’t work in the LCU case: the same rescaling would also be applied to the eigenvalues of $\mathcal{H}$ so that the success probability would stay constant.
In Fig. 9, we have plotted the post-selection probabilities for the physical systems presented in section VI. We considered different regimes for every Hamiltonian, and we have fixed small parameters to reduce $N_{\text{MPO}}$ as much as possible. We have also considered one case for every system with larger parameters to show the sharper decay that is encountered when $\mathcal{H}$ is not adequately rescaled. The equivalent LCU graph doesn’t show any difference by an equal rescaling of the Hamiltonian’s parameters, which confirms our previous point.
We can argue that the MPO block encoding shows an exponential suppression on the asymptotic limit. However, we are still able to get functional probabilities perfectly comparable to the LCU method when considering a finite number of lattice sites and by applying a smart rescaling of $\mathcal{H}$.
The exponential scaling of $p_{\text{success}}$ still represents the biggest limitation of our method. However, we have shown that for a wide range of circumstances, we are still able to find probabilities that are comparable to the LCU block encoding.